The hail comes straight down at a mass rate of m/Δt=0.0690 kg/s and an initial velocity of v0= -15.0 m/s. Suppose that the hail described there bounces off the roof of the car with a velocity of +12.6 m/s. Ignoring the weight of the hailstones, calculate the magnitude of the force exerted by the hail on the roof.
Answer #1

Initial velocity of the hail = V0 = - 15m/s

Velocity of the hail after bouncing off the roof of the car = V1 = 12.6 m/s

Mass flow rate of hail = m/t = 0.069 kg/s

Force exerted by the hail on the roof = F

Force is equal to the rate of change of momentum

F = (0.069)(12.6 - (-15))

F = 1.904 N

Magnitude of the force exerted by the hail on the roof = 1.904 N